OpenStudy (anonymous):

12^6/12^8 what is the simplified form, part a.write your awnser using positive exponents part b what is the value of the expression

OpenStudy (bahrom7893):

12^(-2)

OpenStudy (bahrom7893):

oh wait part a: 1/12^2

OpenStudy (bahrom7893):

And it's value is: 1/144

OpenStudy (bahrom7893):

a^b/a^c = a^(b-c)

OpenStudy (anonymous):

thanks again

OpenStudy (bahrom7893):

1/a^b = a^(-b)